OpenStudy (anonymous):

transform cos 38x to functions of 19x

OpenStudy (anonymous):

have you heard of the trig identy cos(a+b)=cos(a)cos(b)+sin(a)sin(b)? because cos(38x) can be the same as cos(19x+19x) and so a=19x and b=19x

OpenStudy (anonymous):

I Have not heard of that but it makes sense to me. Would that also work to make it a function of 76x?

OpenStudy (anonymous):

\[cos(2x)=1-sin^2(x)\]

OpenStudy (anonymous):

so \[cos(76x)=1-sin^2(38x)\]

OpenStudy (anonymous):

btw calcstud shows how to derive the formula i wrote.

OpenStudy (anonymous):

thank you!